Express the following in terms of x without natural logs.
a) ln(1/e^8x)
b) ln(√ex)

Answer by lwsshak3(11628)

Express the following in terms of x without natural logs.
a) ln(1/e^8x)
=ln1-8xlne
=0-8x
=-8x
..
b) ln(√ex)
=lne^1/2+lnx^1/2
=1/2lne+1/2lnx
=1/2+1/2lnx
=1/2(1+lnx)